OpenStudy (caozeyuan):

can a quadratic form be represented by a non-symmetric matrix? If so, hwo to determine if its postive definite.

OpenStudy (holsteremission):

Sure. Here's an example: \[x^2-xy-y^2=\begin{bmatrix}x\\y\end{bmatrix}^\intercal\begin{bmatrix}1&1\\-2&-1\end{bmatrix}\begin{bmatrix}x\\y\end{bmatrix}\]The polynomial is a quadratic form because it's homogeneous and of order \(2\): \[\begin{align*} f(x,y)=x^2-xy-y^2\implies f(tx,ty)&=(tx)^2-(tx)(ty)-(ty)^2\\[1ex] &=t^2x^2-t^2xy-t^2y^2\\[1ex] &=t^2\left(x^2-xy-y^2\right)\\[1ex] &=t^2f(x,y) \end{align*}\]You can check for positive definiteness by checking the eigenvalues of the matrix. They all have to be positive (and thus real - you'll notice that this particular matrix has complex eigenvalues).

OpenStudy (caozeyuan):

but I was told that eigenvalue check only applies to symmetical ones? so its actually generally true for non symmetrical matrices as well?

OpenStudy (holsteremission):

I suppose the standard definition of positive definiteness only applies to Hermitian matrices (\(\mathbf A=\mathbf A^*\), with \(\mathbf A^*\) the conjugate transpose), or, in the real case, symmetric ones (\(\mathbf A=\mathbf A^\intercal\)). And apparently, the notion of a "quadratic form" depends on the context. I'm interpreting it as referring to a homogenous second order polynomial (see the Wiki page), something of the form \(ax^2+bxy+cy^2\). In linear algebra, it seems, a quadratic form is more specifically the product \(f(\mathbf x)=\mathbf x^\intercal \mathbf{Ax}\), where \(\mathbf A\) is taken to be a symmetric real matrix. Then \(\mathbf A\) is positive definite iff \(f(\mathbf x)>0\) for any \(\mathbf x\). So in this sense, there's no meaning of positive definiteness for the example I used above. You can extend the definition of positive definiteness for asymmetric matrices, though. It involves decomposing \(\mathbf A\) into a linear combination of what are called its symmetric and antisymmetric parts, \[\mathbf A=\color{red}{\mathbf A_S}+\color{blue}{\mathbf A_A}=\color{red}{\frac{\mathbf A+\mathbf A^\intercal}{2}}+\color{blue}{\frac{\mathbf A-\mathbf A^\intercal}{2}}\]Then \(\mathbf A\) can be defined to be positive definite iff \(\mathbf A_S\) is too. \(\mathbf A_S\) is of course symmetric, since \[{\mathbf A_S}^\intercal=\left(\frac{\mathbf A+\mathbf A^\intercal}{2}\right)^\intercal=\frac{\mathbf A^\intercal+\left(\mathbf A^\intercal\right)^\intercal}{2}=\mathbf A_S\]and in fact generates the same quadratic form. So we're really not changing much as far as the definition of positive definiteness goes.